OpenStudy (abbycross167):

Can someone show me how to do this? Write the slope-intercept (y = mx + b) form of the equation from the information given. Slope = -2 and the line goes through the point (-1,3)

OpenStudy (abbycross167):

3=2+-1?

TheSmartOne (thesmartone):

y = mx + b where m is the slope and b is the y-intercept First plug in the slope, which they gave you as -2 in place of m. So first plug in -2 in place of m. What do you get? :)

OpenStudy (abbycross167):

y = -2x + b

TheSmartOne (thesmartone):

Ok, now they also gave us the point (-1,3) We need to plug that in to the equation and solve for b. y = -2x + b (-1,3) is in the form of (x,y) So x = -1 y = 3 Plug that into y = -2x + b, and solve for b :)

TheSmartOne (thesmartone):

Ah, @skullpatrol seems to be a mistake in your answer. You plugged in x = -2. The question states x = -1

OpenStudy (abbycross167):

idk what 'b' could be... and @skullpatrol it's ok

TheSmartOne (thesmartone):

That's what we need to solve for! @abbycross167 :)

TheSmartOne (thesmartone):

So, you plug in x = -1, and y = 3 into y = -2x + b and then we solve for b :)

OpenStudy (abbycross167):

3 = -2(-1) + b b= 2?

OpenStudy (skullpatrol):

3 = -2(-1) + b 3 = 2 + b 1 = b

OpenStudy (abbycross167):

so it'll be 3 = -2+ 1?

TheSmartOne (thesmartone):

The final equation would be in the form of y = mx + b where m is the slope b is the y-intercept we got m = -2 and b = 1 so we finally get y = -2x + 1

OpenStudy (abbycross167):

oh ok, I get it now!!! thank yall very much!!

OpenStudy (skullpatrol):

Thanks for trying :D

TheSmartOne (thesmartone):

Anytime! :D